Suppose the population (age 16 and over) of Italy is 48 million; 4 million are unemployed, and 40 million hold jobs. Indicate the rate of unemployment and the employment/population ratio of New Zealand.

a.

The unemployment rate is 10 percent, and the employment/population ratio is 75 percent.

b.

The unemployment rate is 9.1 percent, and the employment/population ratio is 91.7 percent.

c.

The unemployment rate is 10 percent, and the employment/population ratio is 83.3 percent.

d.

The unemployment rate is 9.1 percent, and the employment/population ratio is 83.3 percent.

During an economic boom,

a.

the actual rate of unemployment will exceed the natural rate of unemployment.

b.

the output of the economy will exceed its long-run potential output.

c.

widespread unemployment will cause inflation to increase.

d.

the actual rate of unemployment will equal the natural rate of unemployment.
Answer #1

1) Unemployment rate = unemployment / labor force = 4*100/(40 + 4) = 9.09%.

Employment population ratio = 40*100/48 = 83.33%.

Option D is correct

2) Option B. Actual unemployment rate is less than its full employment and real GDP is greater than its full

employment level during expansion.